A charge +Q is located at the origin and a negative charge, -9Q, is located at a distance x= 11.0 cm . Where should a third charge q be placed so that the net force on q is zero? Find q 's position on x-axis.
Answer #1

Let it be placed at x cm away from Q charge.

So,

Equating the force on this charge due to both these charges.

\frac{kq}{x^{2}}=\frac{k9q}{(x+0.11)^{2}}

x+0.11=3x

2x=0.11

x=0.055m

or 5.5cm from charge q.(on the left of q charge, because there will be no position in between both cahrges where force is zero.)